# Causal Panel Analysis under Parallel Trends:

### v4

@@ 2.  TWFE and Its Pitfalls
where $a$ and $b$ are the number of lag and lead terms (BJS, 2024). In the social science literature, the typical practice is to exclude $l=0$, which corresponds to the time period immediately before the transition into the treatment phase, and use it as a reference period (Roth, 2022). Conventionally, $\hat{\delta}^{TWFE}_{l}$ is interpreted as an estimate of $\delta_{l}$ or as a meaningful weighted average of pertinent ITE. Meanwhile, $\hat{\delta}^{TWFE}_{b+}$ is viewed as an estimate for the long-term effect.
3.  HTE-Robust Estimators
@@
In this section, we offer a brief overview and comparison of several recently introduced HTE-robust estimators. We use the term “HTE” to refer to ITE that are nonidentical, i.e. $\tau_{i,t}\neq\tau_{j,s}$ for some $i,j,s,t$. HTE-robust estimators are defined as those that produce causally interpretable estimates under their respective identifying assumptions. For a more comprehensive discussion on all estimators mentioned in this section, please refer to the SM.
@@
Table 1 summarizes the estimators we discuss in this paper. The primary difference resides in the mechanics of their estimation strategies: There are methods based on canonical DIDs and methods based on imputation. We refer to the former as DID extensions and the latter as imputation methods. [^8] Liu, Sha and Zhang (2022) use a similar dichotomy to describe these estimators. DID extensions use DTE, estimated from local, $2\times 2$ DIDs between treated and control observations, as building blocks. Imputation methods use ITE, estimated as the difference between an imputed outcome under control and the observed outcome (under treatment), as building blocks. Imputation methods connect to TWFE through the outcome model, which is fit globally on all available controls, that they use to impute counterfactual outcomes. Different strategies also entail different assumptions. Each DID extension, for example, relies on a particular type of PT assumption, whereas imputation methods presuppose a TWFE model for untreated potential outcomes and require a zero mean for the error terms.
Another noteworthy difference lies in the settings in which these estimators are applicable: Some function exclusively in settings with staggered treatment adoption, while others can accommodate scenarios with treatment reversals. [^9] It is possible to use some methods designed for the staggered setting even when there are reversals by using only a subset of the data—in particular, by removing all observations that come after a treatment reversal—but this would identify a different effect than the ATT. In the latter setting, all estimators we discuss also require no anticipation and no or limited carryover effects. Furthermore, the estimators diverge in terms of (1) how they select untreated observations as controls for treated units, (2) how they incorporate pre-treatment or exogenous covariates, and (3) the choice of the reference period. We discuss these details further below and in Sections A.1. and A.4. of the SM.
@@
Finest estimand $\delta_{g,l}$ $\delta_{g,l}$ $\delta_{l}^{vw}$ $\delta_{l}$ $\delta_{l}$ $\tau_{i,t}$ $\tau_{i,t}$
@@
DID extensions are all built from local, $2\times 2$ DID estimates—hence our choice of terminology. The overarching strategy for these estimators is to estimate the DTE, $\delta_{l}$ (or $\delta_{g,l}$ for each cohort $g$ in the staggered setting), for each period since the most recent initiation of treatment, $l$, using one or more valid $2\times 2$ DIDs. By “valid,” we mean that the DID includes (1) a pre-period and a post-period and (2) a treated group and a comparison group. The pre-period is such that all observations in both groups are in control, while the post-period is such that observations from the treated group are in treatment and those from the comparison group are in control. The choice of the comparison group is the primary distinction between estimators in this category. To obtain higher-level averages such as the ATT, we then average over our estimates of $\delta_{l}$ (or $\delta_{g,l}$), typically employing appropriate, convex weights.
@@
We cover three estimators in this category that are appropriate only for the staggered setting. Sun and Abraham (2021) propose an interaction-weighted (IW) estimator, which is a weighted average of $\delta_{g,l}$ estimates obtained from a TWFE regression with cohort dummies fully interacted with indicators of relative time to the onset of treatment. They demonstrate that each resulting estimate of $\delta_{g,l}$ can be characterized as a difference in the change in average outcome from a fixed pre-period $s<g$ to a post-period $l$ periods since $g$ between the treated cohort $g$ and the comparison cohort(s) in some set $\mathcal{C}$. The authors recommend using $\mathcal{C}={\sup_{i}E_{i}}$, which is either the never-treated cohort or, if no such cohort exists, the last-treated cohort. By default, IW uses $l=0$ as the reference period and can accommodate covariates in the TWFE regression.
@@
Employing the same general approach, Callaway and Sant’Anna (2021) propose two estimators, one of which uses never-treated units ($\hat{\delta}^{CS-dr}_{nev}$) and the other not-yet-treated units ($\hat{\delta}^{CS-dr}_{ny}$) as the comparison group. We label these estimators collectively as CSDID. Note that $\hat{\delta}^{CS-dr}_{nev}$ uses the same comparison group as IW when a never-treated cohort exists, [^10] This equivalence holds when there are no missing data. Otherwise, IW from the saturated regression differs from one that directly estimates local DIDs, including the never-treated version of CSDID. These estimates are typically close but can differ substantially, as in Kuipers and Sahn (2023). whereas $\hat{\delta}^{CS-dr}_{ny}$ uses all untreated observations of not only never-treated units but also later adopters as controls for earlier adopters. Besides the choice of comparison cohort, CSDID estimators differ from IW in that they allow users to condition on pre-treatment covariates using both an explicit outcome model and inverse probability weighting (IPW) simultaneously, with at least one needing to be correct for the estimator to be consistent. By default, both IW and CSDID use the period immediately before the treatment’s onset as the reference period for estimating the ATT.
@@
Stacked DID is another related estimator sometimes used to address HTE concerns. As described by Baker, Larcker and Wang (2022), it involves creating separate sub-datasets for each treated cohort by combining data from that cohort (around treatment adoption) and data from the never-treated cohort from the same periods. These cohort-specific datasets are then “stacked” to form a single dataset. An event study regression akin to Model (2) with the addition of sub-dataset specific unit and time dummies is then run. This method uses the same comparison group as IW and the never-treated version of CSDID without covariates, but stacked DID estimates a single DTE for a given relative period rather than separate estimates for each cohort. Essentially, stacked DID is a special case of IW that uses immutable weights selected by OLS. We denote the corresponding estimand $\delta_{l}^{vw}$ to reflect the fact that it is variance-weighted. These weights will be convex but in general are not proportional to cohort sizes. Thus, while stacked DID avoids the “negative weighting” problem and meets our criteria for HTE-robustness, its estimands are not the same as $\delta_{l}$ or the ATT. This, however, is arguably a minor issue related to external validity, as the estimands can be interpreted as average treatment effects for populations with a different distribution over ITE than the treated observations.
@@
In settings with treatment reversals, separate groups of researchers have converged on the same strategy for choosing a comparison group: matching treated and control observations that belong to units with identical treatment histories. IKW (2023) suggest one such estimator, PanelMatch, which begins by constructing a “matched set” for each observation $(i,t)$ such that unit $i$ transitions into treatment at time $t$. This matched set includes units that both (1) are not under treatment at time $t$ and (2) share the same treatment history as $i$ for a fixed number of periods leading up to the treatment onset. For each treated observation $(i,t)$ and for every post-period $t+l-1$ such that unit $i$ is still under treatment, it then estimates a local DID using the same pre-period $s<t$. The treatment “group” comprises solely of $(i,t)$, and the members of the matched set for $(i,t)$ that are still under control during $t+l-1$ serve as the comparison group. To obtain $\delta_{l}$ for a given $l$, it then averages over the corresponding local DID estimates from all treated observations. IKW (2023) propose incorporating covariates by “refining” matched sets and use $l=0$ as the reference period.
[…]
@@
Although DID extensions and imputation methods rely on slightly different identification assumptions, these assumptions usually lead to similar observable implications. Researchers commonly use the presence or absence of pretrends to judge how plausible the PT assumption is. In the classic two-group setting, if there are data from multiple pre-treatment periods, researchers can plot the time series of average outcomes of each group and visually inspect whether they indeed trend together. The intuition is that if PT holds and the outcome trends of the treated and control groups are indeed parallel in pre-treatment periods when $Y(0)$’s are observed for all units, then it is plausible that PT also holds in the post-treatment periods, when $Y(0)$’s are no longer observable for units in the treatment group. Conversely, differential trends in the pre-treatment periods should make us suspicious of PT. In more complex settings or where we wish to control for observed confounders, researchers often use estimates of the dynamic effects before and after the onset of treatment, $\delta_{l}$, to construct so-called “event-study plots” to judge the presence of pretrends. If PT holds, then pre-treatment DTE estimates should be around zero. We provide a more thorough discussion and an example of the event-study plot in the next section when we introduce our procedure.
@@
In general, we believe the credibility of identifying assumptions is more important than the choice of estimator. After all, in the staggered setting, when assumptions hold, IW, CSDID, PanelMatch, DIDM, and the imputation estimator all converge to the same or a similar estimand. However, there are a few reasons to favor imputation estimators. First, imputation estimators can handle complex settings, including those with treatment reversal—which account for over half of the studies in our sample—and can accommodate time-varying covariates, additional fixed effects, and unit- or group-specific time trends commonly seen in social science research. Second, imputation estimators are the most efficient under homoskedastic errors (BJS, 2024). [^12] In Section A.4.2 of the SM, we demonstrate that the imputation estimator is generally more efficient, based on data from our sample. Moreover, by using the average of all pre-treatment periods as the reference point rather than a single pre-period, as in DID extensions, they provide greater power in hypothesis testing for pretrends (see Section A.1.6 of the SM for Monte Carlo evidence). The main drawback of imputation estimators is that their current implementations (either FEct or DID_impute) does not allow for automated adjustment of time-invariant covariates, an advantage offered by CSDID and PanelMatch. Adjusting for pre-treatment characteristics can improve credibility of research, as conditional PT may be more plausible than the unconditional one (Sant’Anna and Zhao, 2020).
4.  Data and Procedure
Next, we assess the robustness of empirical findings from causal panel analyses in political science and compare results obtained using the different methods we have discussed. We will explain our sample selection rules, describe standard practices in the field, and outline our reanalysis approach. Readers can find a more detailed explanation of our sample selection criteria and replication and reanalysis procedure in Sections A.3. and A.4. of the SM.
@@ 4.  Data and Procedure
Our replication sample comprises articles from three leading political science journals, the APSR, AJPS, and JOP, published over a recent seven-year span from 2017 to 2023. We initially include all studies, including both long and short articles, that employ panel data analyses with a binary treatment as a crucial component of their causal argument, resulting in a total of 102 articles. After a careful review of each of these articles, we find that 64 articles use a TWFE model similar to Model (1). We then attempt to replicate the main results of these 64 articles and are successful in 49 cases (76.6%). Though a significant proportion of studies failed to replicate, we note that the success rate is still higher than that of Hainmueller, Mummolo and Xu (2019) at 55% and is comparable to that of Lal et al. (2024), which stands at 67%. We credit this to the new replicability standards set by journals. A detailed explanation of how we select the “main model” is provided in Section A.3. of the SM. Table 2 depicts the distribution of successful replications, along with reasons for replication failures, across the various journals.
Table 2: Sample Selection and Replicability of Qualified Articles
@@
Table 3 presents an overview of the standard practices and settings in the articles that we successfully replicated. The vast majority of studies in our sample (93.9%) use the DID design/method/approach to justify the use of the TWFE model, while the remaining studies advocate for the model’s ability to exploit “within” variations in the data. Out of the 49 articles, nine (18.4%) employ a classic DID design, which includes two-group, two-period designs (3 articles) and multi-period DID designs (6 articles). 13 articles (26.5%) use a staggered (but not classic) DID design, while the remaining 27 articles (55.1%) fall into the “general” category, meaning they allow for treatment reversals. Except for four articles, all studies have a continuous outcome of interest. Most studies adopt cluster-robust SEs or panel-corrected SEs (Beck and Katz, 1995), while eight articles apply bootstrap procedures for estimating uncertainties. A subset of authors explore alternative model specifications by adding lagged dependent variables (8 articles), unit-specific linear time trends (15 articles), and higher-than-unit-level time trends (5 articles). Notably, 30 studies use some type of visual inspection—either average outcomes over time, event-study plots, or both—to evaluate the plausibility of PT. Four studies published in 2023 (33%) employ HTE-robust estimators, compared to none before 2023, indicating rapid adoption of these methods. Of these, two use CSDID, one PanelMatch, and one the imputation estimator. Roughly 40% of the studies in our sample do not include any visual inspection, such as an event-study plot, in the main text or appendices.
@@
“Difference-in-differences” 46 93.9% Cluster-robust SE or PCSE 48 98.0%
“Within” variations 3 6.1% Cluster-bootstrapped procedures 8 16.3%
@@
Classic multi-period DID 6 12.2% Higher-than-Unit-level time trends 5 10.2%
@@
We use data from Grumbach and Sahn (2020) to illustrate our process for replication and reanalysis. The study investigates the influence on coethnic mobilization by minority candidates during US congressional elections. To simplify our analysis, we focus on the impact of the presence of an Asian candidate on the proportion of general election contributions from Asian donors. To begin, we aim to understand the research setting and data structure. We visualize the patterns of treatment and outcome variables using plots, which are shown in the SM. In this application, treatment reversals clearly take place. Some data is missing (due to redistricting), but the issue does not seem to be severe. We record important details such as the number of observations, units, and time periods, the type of variance estimator, and other specifics of the main model (not shown here). Next, we replicate the main finding, employing both the original variance estimator and a cluster-bootstrap procedure.
We then re-estimate the ATT and DTE using estimators discussed in Section 3. For staggered adoption treatment cases, we apply seven estimators: TWFE (with always treated units removed for easier comparisons with other estimators), the imputation estimator (FEct), PanelMatch, DIDM, StackedDID, IW, and CSDID (both not-yet-treated and never-treated versions). For applications with treatment reversals like Grumbach and Sahn (2020), we use the first two estimators only. [^13] When there are treatment reversals, the current release of the DIDM software implements a different estimator than the one described in de Chaisemartin and D’Haultfœuille (2020). In particular, it only estimates the contemporaneous effect at the first onset of treatment as if there are no treatment reversals. Therefore, we only apply DIDM to the staggered cases. The comparison between the TWFE estimate and the other estimates sheds light on whether original findings are sensitive to relaxing the constant treatment effect assumption. Figures 2(a)-(d) show the results from this example. The similarity between estimates for the ATT in (a) suggests that the original finding is robust to the choice of estimators. The event-study plots from HTE-robust estimators in (c) and (d) are broadly consistent with the event-study plot from TWFE in (b).
@@
Note: Reanalysis of data from Grumbach and Sahn (2020). Subfigures (a): Treatment effect or ATT estimates from multiple methods; subfigures (b)-(d): event-study plots using TWFE, PanelMatch, and the imputation estimator (FEct); Subfigure (e)-(f): results from the placebo test and test for carryover effects using the imputation estimator—the blue points in (g) and red points in (h) represent the holdout periods in the respective tests. CIs in all subfigures—excepted for the reported estimate in (a)—are produced by bootstrap percentile methods. The robust confident set for the ATT ($\bar{M}=0.5$) is shown in pink in (e).

### v5

@@ Abstract
where $a$ and $b$ are the number of lag and lead terms (BJS, 2024). In the social science literature, the typical practice is to exclude $l=0$, which corresponds to the time period immediately before the transition into the treatment phase, and use it as a reference period (Roth, 2022). Conventionally, $\hat{\tau}^{TWFE}_{l}$ is interpreted as an estimate of $\tau_{l}$ or as a meaningful weighted average of pertinent ITE. Meanwhile, $\hat{\tau}^{TWFE}_{b+}$ is viewed as an estimate for the long-term effect.
HTE-Robust Estimators
@@
In this section, we offer a brief overview and comparison of several recently introduced HTE-robust estimators. We use the term “HTE” to refer to ITE that are nonidentical, i.e. $\tau_{i,t}\neq\tau_{j,s}$ for some $i,j,s,t$. HTE-robust estimators are defined as those that produce causally interpretable estimates under their respective identifying assumptions. For a more comprehensive discussion on these estimators, please refer to the SM.
@@
Table 1 summarizes the estimators we discuss in this paper. The primary difference resides in the mechanics of their estimation strategies: There are methods based on canonical DIDs and methods based on imputation. We refer to the former as DID extensions and the latter as imputation methods. DID extensions use DTE, estimated from local, $2\times 2$ DIDs between treated and control observations, as building blocks. Imputation methods use ITE, estimated as the difference between an imputed outcome under control and the observed outcome (under treatment), as building blocks. The imputation estimator we use in this paper employs TWFE, fitted with all observations under the control condition, to impute treated counterfactuals. Different strategies also entail different assumptions. Each DID extension, for example, relies on a particular type of PT assumption, whereas imputation methods presuppose a TWFE model for untreated potential outcomes and require a zero mean for the error terms.
Another noteworthy difference lies in the settings in which these estimators are applicable: Some estimators function exclusively in settings with staggered treatment adoption, while others can accommodate scenarios with treatment reversals. In the latter setting, all estimators we discuss also require no anticipation and no or limited carryover effects. Furthermore, the estimators diverge in terms of (1) how they select untreated observations as controls for treated units, (2) how they incorporate pre-treatment or exogenous covariates, and (3) the choice of the reference period. We discuss these details further below and in Section A.1 of the SM.
@@
Finest estimand $\tau_{g,l}$ $\tau_{g,l}$ $\tau_{l}^{vw}$ $\tau_{l}$ $\tau_{l}$ $\tau_{i,t}$ $\tau_{i,t}$
@@
DID extensions are all built from local, $2\times 2$ DID estimates. The overarching strategy for these estimators is to estimate the DTE, $\tau_{l}$ (or $\tau_{g,l}$ for each cohort $g$ in the staggered setting), for each period since the most recent initiation of treatment, $l$, using one or more valid $2\times 2$ DIDs. By “valid,” we mean that the DID includes (1) a pre-period and a post-period and (2) a treated group and a comparison group. The pre-period is such that all observations in both groups are in control, while the post-period is such that observations from the treated group are in treatment and those from the comparison group are in control. The choice of the comparison group is the primary distinction between estimators in this category. To obtain higher-level averages such as the ATT, we then average over our estimates of $\tau_{l}$ (or $\tau_{g,l}$), typically employing appropriate, convex weights.
@@
Three estimators in this category are appropriate only for the staggered setting. Sun and Abraham (2021) propose an interaction-weighted (IW) estimator, which is a weighted average of $\tau_{g,l}$ estimates obtained from a TWFE regression with cohort dummies fully interacted with indicators of relative time to the onset of treatment. They demonstrate that, in a balanced panel, each resulting estimate of $\tau_{g,l}$ can be characterized as a difference in the change in average outcome from a fixed pre-period $s<g$ to a post-period $l$ periods since $g$ between the treated cohort $g$ and the comparison cohort(s) in some set $\mathcal{C}$. The authors recommend using $\mathcal{C}={\sup_{i}E_{i}}$, which is either the never-treated cohort or, if no such cohort exists, the last-treated cohort. By default, IW uses $l=0$ as the reference period and can accommodate covariates in the TWFE regression.
@@
Employing the same general approach, Callaway and Sant’Anna (2021) propose two estimators, one of which uses never-treated units ($\hat{\tau}^{CS}_{nev}$) and the other not-yet-treated units ($\hat{\tau}^{CS}_{ny}$) as the comparison group. We label these estimators collectively as CSDID. Note that $\hat{\tau}^{CS}_{nev}$ uses the same comparison group as IW when a never-treated cohort exists, [^9] This equivalence holds when there are no missing data. Otherwise, IW from the saturated regression differs from one that directly estimates local DIDs, including the never-treated version of CSDID. These estimates are typically close but can differ substantially, as in Kuipers and Sahn (2023). whereas $\hat{\tau}^{CS}_{ny}$ uses all untreated observations of not only never-treated units but also later adopters as controls for earlier adopters. Besides the choice of comparison cohort, CSDID estimators differ from IW in that they allow users to condition on pre-treatment covariates using both an explicit outcome model and inverse probability weighting simultaneously; consistency of the estimators requires at least one of these to be correct. By default, both IW and CSDID use the period immediately before the treatment’s onset as the reference period for estimating the ATT.
@@
Stacked DID is another related estimator sometimes used to address HTE concerns. As described by Baker, Larcker and Wang (2022), it involves creating separate sub-datasets for each treated cohort by combining data from that cohort (around treatment adoption) and data from the never-treated cohort from the same periods. These cohort-specific datasets are then “stacked” to form a single dataset. An event study regression akin to Model (2) with the addition of sub-dataset specific unit and time dummies is then run. This method uses the same comparison group as IW and the never-treated version of CSDID without covariates, but stacked DID estimates a single DTE for a given relative period rather than separate estimates for each cohort. Essentially, stacked DID is a special case of IW that uses immutable weights selected by OLS. We denote the corresponding estimand $\tau_{l}^{vw}$ to reflect the fact that it is variance-weighted. These weights will be convex but in general are not proportional to cohort sizes. Thus, while stacked DID avoids the “negative weighting” problem and meets our criteria for HTE-robustness, its estimands are not the same as $\tau_{l}$ or the ATT. This, however, is arguably a minor issue related to external validity, as the estimands can be interpreted as average treatment effects for populations with a different distribution over ITE than the treated observations.
@@
In settings with treatment reversals, separate groups of researchers have converged on the same strategy for choosing a comparison group: matching treated and control observations that belong to units with identical treatment histories. IKW (2023) suggest one such estimator, PanelMatch, which begins by constructing a “matched set” for each observation $(i,t)$ such that unit $i$ transitions into treatment at time $t$. This matched set includes units that both (1) are not under treatment at time $t$ and (2) share the same treatment history as $i$ for a fixed number of periods leading up to the treatment onset. For each treated observation $(i,t)$ and for every post-period $(t+l-1)$ such that unit $i$ is still under treatment, it then estimates a local DID using the same pre-period $s<t$. The treatment “group” comprises solely of $(i,t)$, and the members of the matched set for $(i,t)$ that are still under control during $t+l-1$ serve as the comparison group. To obtain $\tau_{l}$ for a given $l$, it then averages over the corresponding local DID estimates from all treated observations. IKW (2023) propose incorporating covariates by “refining” matched sets and use $l=0$ as the reference period.
Using a similar strategy, de Chaisemartin and D’Haultfœuille (2020) propose a “multiple DID” estimator, DID_multiple. A notable difference is that they include local DIDs for units leaving the treatment and not just those joining the treatment; when there are no treatment reversals or covariates, DID_multiple is a special case of PanelMatch. The original proposal for DID_multiple also only considers the case where we match on a single period and where $l=1$, but since it has been extented. [^10] The most recent implementation, DIDmultiplegtDYN, allows for estimation of dynamic and long-term effects, bringing DID_multiple closer in functionality to PanelMatch. Consequently, the target estimand is not the ATT but rather an average of the contemporaneous effects of “switching” (i.e., the effect of joining or the negative of the effect of leaving at the time of doing so). In the staggered setting, the PanelMatch estimator aligns with the not-yet-treated version of CSDID (without covariate adjustment). We delve into details on the connections between these three estimators in the SM.
@@
Although DID extensions and imputation methods rely on slightly different identification assumptions, these assumptions usually lead to similar observable implications. Researchers commonly use the presence or absence of pretrends to judge how plausible the PT assumption is. In the classic two-group setting, if there are data from multiple pre-treatment periods, researchers can plot the time series of average outcomes of each group and visually inspect whether they indeed trend together. The intuition is that if PT holds and the average outcome trends of the treated and control groups are indeed parallel in pre-treatment periods when $Y(0)$’s are observed for all units, then it is plausible that PT also holds in the post-treatment periods, when $Y(0)$’s are no longer observable for units in the treatment group. Conversely, differential trends in the pre-treatment periods should make us suspicious of PT. In more complex settings or where we wish to control for observed confounders, researchers often use DTE estimates before and after the onset of treatment, $\tau_{l}$, to construct so-called “event-study plots” to judge the presence of pretrends. If PT holds, then pre-treatment DTE estimates should be around zero. We provide a more thorough discussion and an example of the event-study plot in the next section when we introduce our procedure.
@@
In general, we believe the credibility of identifying assumptions is more important than the choice of estimator. After all, in the staggered setting, when assumptions hold, IW, CSDID, PanelMatch, DID_multiple, and the imputation estimator all converge to the same or a similar estimand. However, there are a few reasons to favor imputation estimators. First, imputation estimators can handle complex settings, including those with treatment reversal—which account for over half of the studies in our sample—and can accommodate time-varying covariates, additional fixed effects, and unit- or group-specific time trends commonly seen in social science research. The imputation estimator connects to TWFE through a shared outcome model, and thus any of the aforementioned modifications to the outcome model can be directly mirrored. DID extensions relates to TWFE through their shared connection to DID in the two-unit, two period setting. Classic DID’s inability to naturally accommodate these complexities limits DID extensions on this front. Just like TWFE, the imputation estimator risks misspecification bias, and adding more redundant terms may significantly increase variance. However, we still consider the added flexibility to be an advantage. Second, imputation estimators are the most efficient under homoskedastic errors (BJS, 2024). [^11] In Section A.3 of the SM, we demonstrate that the imputation estimator tends to yield larger $z$-scores, based on data from our sample. In the majority of cases, the imputation estimator has a smaller standard error, and the difference can be especially dramatic for IW and the never-treated version of CSDID, which often discard the vast majority of untreated observations. Moreover, by using the average of all pre-treatment periods as the reference point rather than a single pre-period, as the default in DID extensions, they provide greater power in hypothesis testing for pretrends. The main drawback of imputation estimators is that their current implementations (either FEct or DID_impute) do not allow for automated adjustment of time-invariant covariates, an advantage offered by CSDID and PanelMatch. Adjusting for pre-treatment characteristics can improve credibility of research, as conditional PT may be more plausible than the unconditional one (Sant’Anna and Zhao, 2020).
Data and Procedure
Next, we assess the robustness of empirical findings from causal panel analyses in political science and compare results obtained using the different methods we have discussed. We will explain our sample selection rules, describe standard practices in the field, and outline our reanalysis approach. Readers can find a more detailed explanation of our sample selection criteria and replication and reanalysis procedure in Section A.2 of the SM.
@@
Our replication sample comprises studies from three leading political science journals, APSR, AJPS, and JOP, published over a recent seven-year span from 2017 to 2023. We initially include all studies, including both long and short articles, that employ panel data analyses with a binary treatment as a key component of their causal argument, resulting in a total of 102 studies. After a careful review, as explained in footnote 1, we find that 64 studies employ a TWFE model similar to Model (1). We then attempt to replicate the main results of these 64 studies and are successful in 49 cases (76.6%). Though a significant proportion of studies failed to replicate, we note that the success rate is still higher than that of Hainmueller, Mummolo and Xu (2019) at 55% and is comparable to that of Lal et al. (2024), which stands at 67%. We credit this to the new replicability standards set by journals. A detailed explanation of how we select the “main model” is provided in Section A.3 of the SM. Table 2 depicts the distribution of successful replications, along with reasons for replication failures, across the various journals.
Table 2: Sample Selection and Replicability of Qualified Studies
@@
Table 3 presents an overview of the standard practices and settings in the studies that we successfully replicated. The majority of studies in our sample (67.3%) use the DID design/method/approach to justify the use of the TWFE model, while the remaining studies advocate for the model’s ability to exploit “within” variations in the data. Out of the 49 studies, nine (18.4%) employ a classic block DID setting, which includes two-group, two-period designs (3 studies) and multi-period block DID designs (6 studies). 13 studies (26.5%) use a staggered DID design, while the remaining 27 studies (55.1%) fall into the “general” category, meaning they allow for treatment reversals. Except for four, all studies feature a continuous outcome of interest. Most use cluster-robust SEs or panel-corrected SEs (Beck and Katz, 1995), and eight studies employ bootstrap procedures for estimating uncertainties. A subset of studies explore alternative model specifications by adding lagged dependent variables (8 studies), unit-specific linear time trends (15 studies), and higher-than-unit-level time trends (1 article). Notably, 32 studies use some type of visual inspection—either average outcomes over time, event-study plots, or both—to evaluate the plausibility of PT. Four studies published in 2023 (33%) employ HTE-robust estimators, compared to none before 2023, indicating rapid adoption of these methods. Of these, two use CSDID, one PanelMatch, and one the imputation estimator. Roughly 40% of the studies in our sample do not include any visual inspection, such as an event-study plot, in the main text or appendices.
@@
“Difference-in-differences” 33 67.3% Cluster-robust SE or PCSE 48 98.0%
“Within” variations 16 32.7% Cluster-bootstrapped procedures 8 16.3%
@@
Multi-period block DID 6 12.2% Higher-than-Unit-level time trends 1 2.0%
@@
We use data from Grumbach and Sahn (2020) to illustrate our process for replication and reanalysis. The authors investigate coethnic mobilization by minority candidates during US congressional elections. To simplify our analysis, we focus on the impact of the presence of an Asian candidate on the proportion of general election contributions from Asian donors. To begin, we aim to understand the research setting and data structure. We visualize the patterns of treatment and outcome variables using plots, which are shown in the SM. In this application, treatment reversals clearly take place. Some data is missing (due to redistricting), but the issue does not seem to be severe. We record important details such as the number of observations, units, and time periods, the type of variance estimator, and other specifics of the main model. Next, we replicate the main finding, employing both the original variance estimator and a cluster-bootstrap procedure.
We then re-estimate the ATT and DTE using estimators discussed in the previous section. For staggered adoption treatment cases, we apply seven estimators: TWFE (with always treated units removed for easier comparisons with other estimators), the imputation estimator (FEct), PanelMatch, DID_multiple, StackedDID, IW, and CSDID (both not-yet-treated and never-treated versions). For applications with treatment reversals like Grumbach and Sahn (2020), we use the first three estimators only. [^12] DIDmultiplegtDYN (De Chaisemartin and d’Haultfoeuille, 2024) implements a different estimator from the one described in de Chaisemartin and D’Haultfœuille (2020). It defines an estimand that accounts for carryover effects. To remain consistent with the imputation framework, we only apply DID_multiple to cases featuring staggered treatment timing. The comparison between the TWFE estimate and the other estimates sheds light on whether original findings are sensitive to relaxing the constant treatment effect assumption. Figures 2(a)-(d) show the results from this example. The similarity between estimates for the ATT in (a) suggests that the original finding is robust to the choice of estimators. The event-study plots from HTE-robust estimators in (c) and (d) are broadly consistent with the event-study plot from TWFE in (b).
@@
Note: Reanalysis of data from Grumbach and Sahn (2020). Subfigures (a): Treatment effect or ATT estimates from multiple methods; subfigures (b)-(d): event-study plots using TWFE, PanelMatch, and the imputation estimator (FEct); Subfigure (e)-(f): results from the placebo test (and robust CS) and test for carryover effects using FEct—the blue points in (e) and red points in (f) represent the holdout periods in the respective tests. In (e), the green and pink bars represent the 95% robust CSs when $\bar{M}=0$ and $\bar{M}=0.5$, respectively. CIs in all subfigures—excepted for the reported estimate in (a)—are produced by bootstrap percentile methods.